Doc. Math. J. DMV 787

Section 19 History of Mathematics

In case of several authors, Invited Speakers are marked with a ^{∗}. Karine Chemla: History of Mathematics in China: A Factor in

World History and a Source for New Questions . . . . . . . . . . . . . . . . . . . . III 789 Joseph W. Dauben: Marx, Mao and Mathematics: The Politics of

Infinitesimals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 799 Jeremy J Gray: The Riemann-Roch Theorem and Geometry,

1854-1914 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 811

Documenta Mathematica · Extra Volume ICM 1998 · III ·

788

Documenta Mathematica · Extra Volume ICM 1998 · III ·